OpenStudy (anonymous):

Can anyone help me solve these 2 Quadratic equations ? x^2 – 36 = 0 and x^2 = 8x – 12

OpenStudy (anonymous):

1)x=+6 or x=-6

OpenStudy (akshay_budhkar):

x^2=36 X=+6,-6

OpenStudy (anonymous):

2)x^2-8+12=0 x^2-6x-2x+12=0 x(x-6)-2(x-6)=0 (x-2)(x-6)=0 x=2 x=6

OpenStudy (anonymous):

I think its x^2 - 8 - 12 Not +12

OpenStudy (anonymous):

no it is x^2-8x+12

OpenStudy (anonymous):

1)x^2=36 x=+6,-6 2)^2-8+12=0 x^2-6x-2x+12=0 x(x-6)-2(x-6)=0 (x-2)(x-6)=0 x=2 x=6

OpenStudy (anonymous):

oh ok

OpenStudy (dumbcow):

they added 12 on both sides to cancel it on the right side and move it to left side

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

any other question

OpenStudy (anonymous):

No i am good for now thanks! :) you are all my Heros!!!!! :)

OpenStudy (anonymous):

x^2 – 36 = 0 x^2=36 x=6 or -6